OpenStudy (anonymous):

A 1000 kg aircraft going 25 m/s collides with a 1500 kg aircraft that is parked and they stick together after the collision and are going 10 m/s after the collision. If they skid for 5.5seconds before stopping, how far did they skid? Hint: Are the aircraft moving at a constant velocity after the collision or do they experience an acceleration?

OpenStudy (anonymous):

so, after the collision we have a big mass of metal with a speed of 10m/s, if it skids and then stops there should be an acceleration (in this case it should point against the velocity), so the starting velocity is \(v_0=10m/s\) and, if we consider a constant acceleration we can use this equations for velocity and position: \(v=v_0+at\) \((1)\) \(x=v_0t+\frac{1}{2}at^2\) \((2)\) to find when and where the planes stop we must set \(v=0\), and solve for \(t\) and \(x\) Give it a try!

OpenStudy (anonymous):

so if i'm solving for t and setting v to 0. T = -10/a

OpenStudy (anonymous):

yes... and you can plug that in eq(2) replacing \(t\) by \(\frac{-10}{a}\) and solving that for a :)

OpenStudy (anonymous):

is the answer 75 m o-o

OpenStudy (anonymous):

i get x=27.5m sorry i forgot the time was given in the problem! we need to use t=5.5 sec so a=-1.81m/s^2 please check your calculations

OpenStudy (anonymous):

Ok i got the same answer as you ! thankyou !! :D

OpenStudy (anonymous):

great! good work! :)